# INSIGHT: End-to-End Neuro-Symbolic Visual Reinforcement Learning with Language Explanations

### v1

@@ C.2 Analysis of Coordinate Accuracy
Tab. 3 presents a quantitative assessment of the coordinate prediction accuracy achieved by our method. In Tab. A7, we document the experimental results for the additional four tasks. Notably, the F-MAE on Qbert, Seaquest, and MsPacman post-training markedly surpasses the pre-training figures. This improvement suggests that end-to-end fine-tuning significantly enhances the strategic relevance of object coordinate accuracy.
Table A7: MAE and F-MAE for coordinate prediction for Qbert, Seaquest, Breakout and MsPacman. Numbers are multiplied by 100.
@@
Figure A1: Performance of INSIGHT and it variants for Qbert, Seaquest, Breakout and MsPacman.
@@
Figure A2: Effect of regularization coefficient $\lambda_{\text{reg}}$, the width of the EQL actor, the number of hidden layers in the EQL actor, and the weight of $\mathcal{L}_{\text{CNN}}$.
@@ C.3 Extended Ablation Study Details
This subsection extends the ablation studies detailed in Fig. 3 and Fig. 4 to encompass all tasks, as referenced in Fig. A1-Fig. 1(i). Notably, the comprehensive ablation analysis presented in Fig. A1 demonstrates performance declines in most tasks when methods are modified. Furthermore, the consistency observed across various hyperparameters, as shown in Fig. 1(a)-Fig. 1(i), corroborates the findings discussed in Sec. 4.2.
Table A8: Performance of the proposed INSIGHT, Neural and CleanRL (Huang et al., 2022)) on nine Atari tasks. INSIGHT matches the performance of Neural for all nine tasks and outperforms all NS-RL baselines. Our neural baseline is contestable for the CleanRL baseline in most cases.
@@ C.4 Additional Baseline
In this section, we introduce an additional baseline, CleanRL, to evaluate the performance of INSIGHT. CleanRL is an open-source implementation derived from the CleanRL library (Huang et al., 2022). The sole modification involves adjusting the CNN’s output dimension from 512 to 2048 or higher to facilitate a fair comparison with our framework. It is noteworthy that our neural baseline exhibits competitive performance relative to CleanRL. As shown in Tab. A8, INSIGHT achieves superior returns in six tasks, unequivocally demonstrating its capability to outperform existing neural policies in online environments.
@@ C.5 Visualized Coordinate Tracking
In this section, we expand upon the advantages of end-to-end training for coordinate prediction, initially presented in Tab. 3, by offering a more intuitive visualization. Fig. A3 illustrates the object coordinates before and after the training process. For instance, in Figs. 2(b) and 2(a), we note a marked reduction in the pixel coordinate shift of the ball post-training. Similarly, the laser emitted by the agent in Figs. 2(d) and 2(c) and the car on the road in Figs. 2(f) and 2(e) exhibit significantly diminished pixel coordinate shifts following training. Moreover, post-training observations in Figs. 2(h) and 2(g) reveal the successful detection of enemy coordinates at the periphery, a detail that was previously unattainable. Lastly, in Figs. 2(j) and 2(i), the model demonstrates its ability to consistently track the blue enemy post-training, overcoming the initial limitation of losing track of the object.

### v2

@@ C.2 Analysis of Coordinate Accuracy
End-to-End Fine-Tuning Markedly Enhances Coordinate Prediction Accuracy in Policy-Relevant Tasks Tab. 4 presents a quantitative assessment of the coordinate prediction accuracy achieved by our method. In Tab. A7, we document the experimental results for the additional four tasks. Notably, the F-MAE on Qbert, Seaquest, and MsPacman post-training markedly surpasses the pre-training figures. This suggests that end-to-end fine-tuning significantly improves accuracy in policy-relevant object coordinates.
Table A7: INSIGHT demonstrates robustness to hyper-parameters. MAE and F-MAE are used to measure coordinate prediction accuracy for Qbert, Seaquest, Breakout, and MsPacman, with all numbers scaled up by a factor of 100.
@@
Figure A1: Each component of INSIGHT is critical for overall performance. Detailed performance analyses of INSIGHT and its variants across four tasks are presented. The findings in these four environments align with those in the other five environments.
@@
Figure A2: INSIGHT demonstrates robustness to hyper-parameters. Examining the effects of the regularization coefficient $\lambda_{\text{reg}}$, the width and number of hidden layers of the EQL actor, and the weight of $\mathcal{L}_{\text{CNN}}$ on performance. INSIGHT shows substantial robustness to variations in these hyper-parameters.
@@ C.3 Extended Ablation Study Details
Generalizability of Ablation Experiment Conclusions Across Tasks This subsection extends the ablation studies detailed in Fig. 3 and Fig. 4 to encompass all tasks, as referenced in Fig. A1-Fig. 2(i). Notably, the comprehensive ablation analysis presented in Fig. A1 demonstrates performance declines in most tasks when methods are modified. Furthermore, the consistency observed across various hyperparameters, as shown in Fig. 2(a)-Fig. 2(i), corroborates the findings discussed in Sec. 4.2.
Table A8: INSIGHT outperforms CleanRL. Evaluating the performance of the proposed INSIGHT, Neural, and CleanRL (Huang et al., 2022) across nine Atari tasks. INSIGHT matches Neural’s performance on all tasks and surpasses all NS-RL baselines. In most instances, our neural baseline competes closely with the CleanRL baseline.
@@ C.4 Additional Baseline
INSIGHT Outperforms CleanRL In this section, we introduce an additional baseline, CleanRL, to evaluate the performance of INSIGHT. CleanRL is an open-source implementation derived from the CleanRL library (Huang et al., 2022). The sole modification involves adjusting the CNN’s output dimension from 512 to 2048 or higher to facilitate a fair comparison with our framework. It is noteworthy that our neural baseline exhibits competitive performance relative to CleanRL. As shown in Tab. A8, INSIGHT achieves superior returns in six tasks, unequivocally demonstrating its capability to outperform existing neural policies in online environments.
@@ C.5 Visualized Coordinate Tracking
In this section, we expand upon the advantages of end-to-end training for coordinate prediction, initially presented in Tab. 4, by offering a more intuitive visualization. Fig. A3 illustrates the object coordinates before and after the training process. For instance, in Figs. 3(b) and 3(a), we note a marked reduction in the pixel coordinate shift of the ball post-training. Similarly, the laser emitted by the agent in Figs. 3(d) and 3(c) and the car on the road in Figs. 3(f) and 3(e) exhibit significantly diminished pixel coordinate shifts following training. Moreover, post-training observations in Figs. 3(h) and 3(g) reveal the successful detection of enemy coordinates at the periphery, a detail that was previously unattainable. Lastly, in Figs. 3(j) and 3(i), the model demonstrates its ability to consistently track the blue enemy post-training, overcoming the initial limitation of losing track of the object.
C.6 Accuracy Evaluation of the FastSAM and DeAot Method
Our Approach Competes with MOC Solely Through Zero-Shot Generalization To evaluate the reliability of our dataset generation method, we further evaluate the accuracy of bounding boxes in the OCAtari dataset  (Delfosse et al., 2023a). The table below shows the F-score from the Pong, SpaceInvaders and MsPacman environments within the OCAtari dataset, compared with the results of SPACE+MOC, as reported in  (Delfosse et al., 2023d):
Table A9: Our method competes with MOC solely through zero-shot generalization. The F-score evaluation of the FastSAM and DeAot method on OCAtari.
Env INSIGHT SPACE+MOC(w/o OC) SPACE+MOC
Pong 97.5 91.5 87.4
SpaceInvaders 86.5 85.1 85.2
MsPacman 46.5 88.6 90.5
Table A10: Our method achieves high recall solely through zero-shot generalization. The precision and recall evaluation of the FastSAM and DeAot method on OCAtari.
Env Precision Recall
Pong 96.3 98.7
SpaceInvaders 88.5 84.5
MsPacman 30.5 97.8
Note that INSIGHT achieves an impressive F-score of 97.5% on Pong, significantly outperforming SPACE+MOC. For SpaceInvaders, we set IOU to 1 to segment as many different objects as possible. However, MsPacman’s F-score is lower than that achieved with the MOC-based method.
Our Method Achieves High Recall Solely Through Zero-Shot Generalization To further analyze this, we provide the precision and recall scores in Table A10. In the MsPacman environment, the very low precision and nearly perfect recall indicate that while all objects were segmented, many irrelevant objects not in the OCAtari dataset were also included. Given our goal during dataset generation is to segment as many complete objects as possible for downstream symbolic policy, higher precision may be more appropriate than F-score in RL scenarios.
C.7 Additional Computation Time Information
The average training time of each component on 9 Atari environments using NVidia RTX 3090 GPU is as follows:
Table A11: Training time of each component on 9 Atari environments.
Pretraining the Visual Perception Policy learning Policy Learning w/o Finetune
Time (ms) $2.6\pm 0.3$ $7.3\pm 1.1$ $6.2\pm 0.8$
Note that the end-to-end training time of the visual module only accounts for about 1/10 of the total pipeline, but it greatly improves the return and prediction accuracy, which reflects the high profitability.
Overall Computation Time Increases with the Number of Objects  In Tab. A12, we present the computation times of FastSAM, SAM, and DeAot with varying numbers of objects:
Table A12: Overall computation time increases with the number of objects. The computation times (s) of FastSAM, SAM, and DeAot with varying numbers of objects.
Num FastSAM SAM DeAot
5 $0.07$ $4.21$ $0.01$
10 $0.07$ $4.57$ $0.02$
20 $0.07$ $5.21$ $0.02$
50 $0.08$ $5.83$ $0.03$
100 $0.08$ $5.84$ $0.07$
The segmentation speed of FastSAM remains relatively constant despite varying object counts, whereas DeAot’s speed significantly decreases as more objects are introduced. This suggests an increase in overall computational overhead proportional to the number of objects.
The average inference time of each component on 9 Atari environments using NVidia RTX 3090 GPU is as follows
Table A13: Inference time of each component on 9 Atari environments.
CNN EQL
Time (ms) $0.3$ $1.7$
Note that we have adopted a simple CNN as the tracking module, which significantly enhances the inference speed of INSIGHT.
C.8 Dataset segmentation results
To visually demonstrate our segmentation accuracy, we took five frames from each Atari game, and summarized them in Fig. A4:
(a) Pong.
(b) BeamRider.
(c) SpaceInvaders.
(d) Seaquest.
(e) MsPacman.
Figure A4: FastSAM and DeAot Extract Segmentation Results from Five Frames per Game, Sampled Every Five Frames.
Note that our segmentation data set completes accurate segmentation of objects in most cases.